OpenStudy (anonymous):

31.03 is what percent of 53.5? ? %

OpenStudy (superflytyguy):

Just divide: 31.03 / 53.5 = 0.58

OpenStudy (superflytyguy):

Whenever you need to find the percent divide smaller number by bigger number. For example 1/2 = 0.5 or 50%

OpenStudy (superflytyguy):

So your answer is 58%

sammixboo (sammixboo):

Remember what I said earlier? \(\large \rm \frac{part}{whole} ~=~\frac{percent}{100}\) 31.03 is our part 53.5 is our whole. Plug them in! :) \(\large \rm \frac{31.03}{53.5} ~=~\frac{percent}{100}\) First multiply 31.03 by 100, so what is 31.03 x 100?

OpenStudy (anonymous):

3.103

sammixboo (sammixboo):

Ah nope! It is 3103! Now take 3103 and divide it by 53.5 So what is 3103/53.5? Use a calculator if needed

OpenStudy (anonymous):

58

sammixboo (sammixboo):

And that is your answer!

OpenStudy (anonymous):

thanks again :)

sammixboo (sammixboo):

No problem!

OpenStudy (anonymous):

who wants the medal

sammixboo (sammixboo):

@SuperflyTyGuy can have it!

sammixboo (sammixboo):

He helped you first =)

OpenStudy (superflytyguy):

Thanx @sammixboo

sammixboo (sammixboo):

No problem!

OpenStudy (anonymous):

:) ok